6. -1 points OSColPhysAP2016 6.2.WA.022.Tutorial. My Notes Ask Your Teacher 0 A jet fighter flying at 300 m/s (just below the speed of sound) makes a turn of radius 1.90 km. (a) What is its centripetal acceleration in gs? (b) Suppose the pilot makes an emergency turn to avoid an approaching missile, subjecting himself to a centripetal acceleration of 10 gs, while flying at 520 m/s (supersonic). What is the radius, in km, of his turn? (This must be short-lived because fighter planes can only briefly endure such large accelerations without serious damage, and the pilot will soon black out at 10 gs.) km Additional Materials Tutorial
Answer #1

(a) What is its centripetal acceleration in gs?

ac = V^2/(r*g)

= (300)^2/(1.9*10^3*9.8) = 4.83 g

(b) Suppose the pilot makes an emergency turn to avoid an approaching missile, subjecting himself to a centripetal acceleration of 10 gs, while flying at 520 m/s (supersonic). What is the radius of his turn? (This must be short-lived because fighter planes can only briefly endure such large accelerations without serious damage, and the pilot will soon black out at 10 gs.)

10g = V^2/r

98r = 520^2

r = 520^2/98 = 2,759 m = 2.76 km